OpenStudy (anonymous):

find the solutions of the equation 1/2x^2 - x + 5 = 0

OpenStudy (anonymous):

Multiply by 2 both the sides fist: \[x^2 - 2x + 10 = 0\] Now use the quadratic formula to find x..

OpenStudy (lgbasallote):

\[\frac{1}{2x^2} - x + 5 = 0\] multiply everything by 2x^2 \[1 - 2x^3 + 5x^2 = 0\] \[-2x^3 + 5x^2 - 1 = 0\] factor it out

OpenStudy (lgbasallote):

lol jk..that's not the question

OpenStudy (anonymous):

this one really confuses me because my choices have i in them

OpenStudy (anonymous):

\[D = b^2 - 4ac\] D = 4 - 40 = -36 \[x = \frac{-b \pm \sqrt{D}}{2a}\] So, \[x = \frac{2 \pm 6i}{2}\] So, \[\color{green}{x = 1 + 3i} \qquad Or \qquad \color{blue}{x = 1 - 3i}\]

OpenStudy (anonymous):

\[\large \color{red}{\sqrt{-36} = 6i}\]